OpenStudy (anonymous):

Find the equation of a line passing through (-2, 5) and (3, 4). Enter the equation in the form a x + b y = c

OpenStudy (noseboy908):

are you familiar with point-slope form?

OpenStudy (anonymous):

Somewhat.

OpenStudy (noseboy908):

ok, what is it?

OpenStudy (anonymous):

y-y1 = m(x-x1)

OpenStudy (noseboy908):

correct, now, from your two points, what's your slope?

OpenStudy (anonymous):

-1/5

OpenStudy (noseboy908):

yep, so fill in your given point and slope

OpenStudy (anonymous):

I got y = -1/5x + 3. How do I put that into the form they want it in?

OpenStudy (anonymous):

@austinL

OpenStudy (austinl):

Just add \(\dfrac{x}{5}\) to both sides to get, \(\dfrac{x}{5}+y=3\)

OpenStudy (anonymous):

Thanks! :))

OpenStudy (austinl):

No problemo!

OpenStudy (noseboy908):

any other questions?